(a). write the value of the discriminant of each equation. (b). Then use it to decide how many different real number roots the equation has.(do not solve.)
3x to the second power - x - 1 = 0

Answer by fractalier(6550)

For 3x^2 - x - 1 = 0
the discriminant b^2 - 4ac =
1 + 12 = 13
which means there are two conjugate real roots